OpenStudy (osanseviero):

Fast questions (2) about factorization...

OpenStudy (osanseviero):

\[(3x ^{2}-x)\]

OpenStudy (anonymous):

Both terms have an x, so you can take an x out and divide em by x. 3x² - x = x(3x - 1)

OpenStudy (osanseviero):

This is a rectangle...you know the area is 70. The long is 3 times the width -1. Which is the width and which is the long? (you can say width is x, long is 3x-1...and x(3x-1)=70...but what comes then?

OpenStudy (anonymous):

You need to solve for x?

OpenStudy (osanseviero):

Yep

OpenStudy (anonymous):

x(3x - 1) = 70 Multiply out. 3x² - x = 70 Subtract 70. 3x² - x - 70 = 0 See where I'm going? Or do you want me to continue?

OpenStudy (osanseviero):

So now I just need to find two numbers than multiplied are -210 and added -1?

OpenStudy (anonymous):

I think you should just use Bhaskara, it's easier in my opinion

OpenStudy (anonymous):

I like tic tac toe. http://www.slideshare.net/qvamath/tic-tac-toe-factoring 3x² | -70 | -210x² 3x | -5 | -15x x | -14| -14x (3x - 14)(x - 5) = 0 x = 5, 14/3 by zero product property.